OpenStudy (anonymous):

please, i confused bit Evaluate :limx....(-1) x^2+2x+1/x+1

OpenStudy (lgbasallote):

\[\large \lim_{x \rightarrow -1} \frac{x^2 + 2x + 1}{x+1} \implies \frac{(x+1)(x+1)}{x+1} \implies \frac{(x+1)\cancel{(x+1)}}{\cancel{x+1}} \implies x+1\] substitute x now

OpenStudy (lgbasallote):

do you get it?

OpenStudy (anonymous):

yeah this is what thouguht it about

OpenStudy (lgbasallote):

nice. you were on the right track then :D

OpenStudy (lgbasallote):

so can you find the answer now or do you need more help?

OpenStudy (anonymous):

nope thankz i already have the answer but i was a bit doubt